# A Bayesian Nonparametric Stochastic Block Model for Directed Acyclic Graphs

### v2

@@ 3.2 SNA citation network
We first look at the citation network analysed by Lee and Wilkinson (2018). It contains 1118 citations (edges) between 135 articles (vertices) which are related to social network analysis (SNA). We shall call this data the SNA citation network hereafter.
The MCMC sampler for the DAG-SBM outlined in the online Appendix C was applied, with 20000 iterations obtained after a burn-in period of 100000 and thinning of 2000 i.e. keeping 1 sample every 2000 iterations, with $a\sim\text{Gamma}(1,0.01)$, $b\sim\text{Gamma}(1,0.01)$. Such settings were used three times, once assuming the infinite regime $(r=0)$ with $\theta+\alpha\sim\text{Gamma}(1,0.01)$, once assuming the finite regime $(r=1)$ with $\gamma\sim\text{Gamma}(1,0.01)$ and $k\sim$ truncated negative binomial$(1,0.01)$, and once for model selection with $\mathbb{P}(r=1)=0.2$ i.e. $\mathbb{P}(r=0)=0.8$. All three runs were performed on a Linux machine with Intel Core i7-7700 Processor (3.6GHz), and took 0.0089, 0.0077, and 0.0083 seconds per iteration, respectively.
Figures 4 to 9 show some key inference results, except for the trace plots of the parameters, which are in the online Appendix D.2. Of more importance is the posterior density (or mass function in the case of $K_{n}$ and $k$) in Figures 4 and 5. On one hand, in the panels for $k$ and $\gamma$, the infinite regime is naturally missing, while the results by the finite regime and model selection coincide as expected. Similarly, in the panels for $\theta$ and $\alpha$, the finite regime is naturally missing, while the results by the infinite regime and model selection coincide as expected. On the other hand, the panels for $K_{n}$, $a$ and $b$ illustrate that the posterior density from model selection is essentially a weighted average of that of the two regimes.
The departure of the posterior densities of $\alpha$ and $\gamma$ from 0 suggests that either the infinite regime or the finite regime is preferred to their shared boundary i.e. the Dirichlet process when $\alpha=\gamma=0$. Between the two regimes, with $\mathbb{P}(r=1)=0.2$ resulting in $\mathbb{P}(r=1|\mathbf{Y})=0.5474$, the Bayes factor $B_{10}=\displaystyle\frac{\mathbb{P}(r=1|\mathbf{Y})}{\mathbb{P}(r=0|\mathbf{Y})}\left/\frac{\mathbb{P}(r=1)}{\mathbb{P}(r=0)}\right.=4.8368$, suggesting a slight preference for the finite regime, for this citation network.
Figure 4: Posterior histogram of $K_{n}$ and $k$, for the finite (solid) and infinite (stripe) regimes of the DAG-SBM, and model selection (crosshatch), for the SNA citation network.
@@
Next, we look at the posterior of $\boldsymbol{\sigma}$, or equivalently the positions of the vertices in the topological ordering, $\boldsymbol{\phi}$. The mixing of the MCMC is sufficiently good and the posterior density of each component of $\boldsymbol{\phi}$ is plotted as a row in Figure 7, with the rows themselves in an arbitrary topological ordering. As the citations i.e. the edges in the DAG in general go from more recent works to older ones, the top (bottom) rows, which correspond to topologically earlier (later) vertices, are indeed the more recent (older) articles in the network. The coloured part of each row represents the support of the posterior of the corresponding article, and in general is narrower towards the bottom. An interpretation is that chronologically older works are likely to cite between themselves and get cited by later ones, thus limiting their positions in $\boldsymbol{\sigma}$.
Figure 7: The posterior density of the positions of the vertices in $\boldsymbol{\sigma}$ for the infinite (left) and finite (right) regime of the DAG-SBM, for the SNA citation network. Each row of colours is the posterior density of a component of $\boldsymbol{\phi}$. The black dots are the mean positions.
Figure 9: Network diagram according to the point estimate of the allocation vector, $\hat{\mathbf{Z}}_{n}$, for the infinite (left) and finite (right) regime of the DAG-SBM, for the SNA citation network.
To provide a posterior point estimate of $\mathbf{Z}_{n}$, we follow the clustering approach introduced in (Meilă, 2007) and further discussed in (Wade and Ghahramani, 2018). Specifically, the point estimate, denoted by $\hat{\mathbf{Z}}_{n}$, is obtained using a decision theoretic approach, by minimising with respect to the posterior distribution a loss function on the space of allocation vectors,
For the loss function $L(\mathbf{Z}_{n},\tilde{\mathbf{Z}}_{n})$, (Meilă, 2007) chose the Variation of Information (VI), defined as
$\text{VI}(\mathbf{Z}_{n},\tilde{\mathbf{Z}}_{n})=\sum_{i=1}^{K_{n}}\frac{n_{i+}}{n}\log\left(\frac{n_{i+}}{n}\right)+\sum_{j=1}^{\tilde{K}_{n}}\frac{n_{+j}}{n}\log\left(\frac{n_{+j}}{n}\right)-2\sum_{i=1}^{K_{n}}\sum_{j=1}^{\tilde{K}_{n}}\frac{n_{ij}}{n}\log\left(\frac{n_{ij}}{n}\right)$
where the counts variables are defined as $n_{ij}=\sum_{p=1}^{n}\mathbb{I}(Z_{p}=i,\tilde{Z}_{p}=j)$, $n_{i+}=\sum_{j=1}^{\tilde{K}_{n}}n_{ij}$, and $n_{+j}=\sum_{j=1}^{K_{n}}n_{ij}$. The loss function $L(\mathbf{Z}_{n},\tilde{\mathbf{Z}}_{n})$ can be seen as a distance between $\mathbf{Z}_{n}$ and $\tilde{\mathbf{Z}}_{n}$, which can be computed even if $K_{n}\neq\tilde{K}_{n}$ i.e. the numbers of groups implied by $\mathbf{Z}_{n}$ and $\tilde{\mathbf{Z}}_{n}$ are different. Using the SALSO algorithm (Dahl et al., 2022) to minimise the loss function and obtaining the point estimate $\hat{\mathbf{Z}}_{n}$ for the SNA citation network, we plot the co-clustering and adjacency matrices in Figure 8, and the network diagram in Figure 9.
@@ 3.3 Statistics citation network
We apply the model to a second citation network. The original data analysed by (Ji and Jin, 2016) contains 5722 citations between 3248 articles in the top statistics journals. We only consider the largest connected component, and remove one of the two edges in every pair of cyclic citations (there were only 9 pairs of these). Upon the data cleaning, we arrive at a citation network that is a DAG and contains 5563 citations between 2248 articles. We refer to this data as the statistics citation network hereafter.
The MCMC sampler was applied with $1.5\times 10^{4}$ iterations obtained after a burn-in period of $5.5\times 10^{4}$ and thinning of 10, with the same priors for $a$, $b$, $\alpha$, $\theta$, $k$ and $\gamma$ as those in Section 3.2. For model selection, even with $\mathbb{P}(r=0)=0.99$, the whole chain stays in the finite regime i.e. $\mathbb{P}(r=1|\mathbf{Y})=1$. Therefore, we shall report the results under the finite regime only, as it is heavily preferred to the infinite counterpart for this network. The trace plots and posterior densities are provided in the online Appendix D.3. The posterior of $K_{n}$ ranges from 92 to 124, which seems reasonable with around 30 articles per cluster.
@@
Table 2: Titles of top articles according to topological ordering for selected groups
@@
Table 3: Titles of top 20 articles according to topological ordering
@@
Also, we report several groups of top articles according to median topological ordering for some large groups in Table 2. Specifically, for a selection of blocks, we report the top 5 articles in topological order within the block. From the articles’ titles, groups seem to roughly correspond to various topics in statistics. For example, group 18 seems to cover false discovery rates, group 45 design and randomisation, group 2 wavelets, group 71 missing/perturbed data, group 25 function linear regression, and group 65 spatial processes. However, note that the reported group memberships are just point estimates obtained with the MCMC draws, and therefore there might be some possibly large posterior uncertainty in the topological ordering for some of these articles.
Lastly, we report in Table 3 some top articles according the median topological ordering, this time without the group memberships. In general, these articles have high number of citations. However, we also notice that some of these articles might not have a very high number of citations. This is due to the fact that these are articles are cited by some other influential articles towards the end of the topological ordering.
@@ 4 Discussion
In this article, we proposed a Bayesian nonparametric SBM for DAGs. Specifically, by conditioning on a latent topological ordering, the likelihood of the data (which is composed of directed edges) becomes that of an undirected graph, i.e. that of an upper diagonal adjecency matrix. The topological ordering is treated as an unknown parameter, endowed with a prior and inferred a posteriori within the MCMC sampler, using a modified Leap-and-Shift proposal. Also, the use of the PY process prior for the allocation vector $\mathbf{Z}_{n}$ allows the model to infer the number of groups $K_{n}$ from the data. Moreover, a model selection step for the two regimes of the PY process can be included within the MCMC sampler.
The model can be generalized in different ways. For example, it can be extended by introducing covariate information, such as the keywords of each document or its publication year. This could be achieved by modelling the degree correction factors with a covariate dependent distribution. Also, in terms of parametrisation, the two regimes of the PY process could be unified so that $\gamma$ and $\alpha$ become one parameter that can take a value between $\infty$- ∞ and $1$. Its posterior density will directly imply which regime is preferred. The main obstacle to overcome here would be the sampling from the non-standard joint parameter space of $\theta$ and $\alpha$ across the two regimes.
Another issue to be resolved is the inference of $k$, which is naturally highly correlated with $K_{n}$, under the finite regime. This is apparent in the parameter trace plots (in online Appendix D.3) for the statistics citation network, while for the SNA citation network the model selection improves the mixing of $k$ in the MCMC. Ideally $k$ is integrated out, but the computations required mean that this is feasible only under certain special cases. Such issue with $k$ remains to be resolved.
@@
Background materials of SBM and Pitman-Yor process, likelihood derivations, MCMC algorithms, and additional plots (appendices.pdf).
@@
R-package “dagsbm” containing code to run the MCMC sampler for the DAG-SBM, and to simulate graphs from the model. The package also contains the dataset of SNA citation network presented in the article. (dagsbm_0.0.7.tar.gz, GNU zipped tar file)
@@
Script for running the MCMC sampler using the R package. (sim_dc.R)
@@ Results:
Pre-generated results from running sim_dc.R, for both the README and reproducibility checks. (results/*sim_dc_alpha=0.00_theta=1.00_n=250*.rds)
@@
Instructions for reproducing the results in the paper using the R package and code. (README.Rmd)

### v3

@@ 3.2 SNA citation network
Next, we look at the citation network analysed by Lee and Wilkinson (2018). It contains 1118 citations (edges) between 135 articles (vertices) which are related to social network analysis (SNA). We shall call this data the SNA citation network hereafter.
The MCMC sampler for the DAG-SBM outlined in the online Appendix C was applied, with 20000 iterations obtained after a burn-in period of 100000 and thinning of 2000, i.e. keeping 1 sample every 2000 iterations, with $a\sim\text{Gamma}(1,0.01)$, $b\sim\text{Gamma}(1,0.01)$. Such settings were used three times, once assuming the infinite regime $(r=0)$ with $\theta+\alpha\sim\text{Gamma}(1,0.01)$, once assuming the finite regime $(r=1)$ with $\gamma\sim\text{Gamma}(1,0.01)$ and $k\sim$ truncated negative binomial(1,0.01)(1,0.01), and once for model selection with $\mathbb{P}(r=1)=0.2$, i.e. $\mathbb{P}(r=0)=0.8$. All three runs were performed on a Linux machine with Intel Core i7-7700 Processor (3.6GHz), and took 0.0089, 0.0077, and 0.0083 seconds per iteration, respectively.
Figures 4 to 8 show some key inference results, except for the trace plots of the parameters, which are in the online Appendix D.2. Of more importance is the posterior density (or mass function in the case of $K_{n}$ and $k$) in Figures 4 and 5. On one hand, in the panels for $k$ and $\gamma$, the infinite regime is naturally missing, while the results for the finite regime and model selection coincide as expected. Similarly, in the panels for $\theta$ and $\alpha$, the finite regime is naturally missing, while the results for the infinite regime and model selection coincide as expected. On the other hand, the panels for $K_{n}$, $a$ and $b$ illustrate that the posterior density from model selection is essentially a weighted average of that of the two regimes.
The departure of the posterior densities of $\alpha$ and $\gamma$ from 0 suggests that either the infinite regime or the finite regime is preferred to their shared boundary, i.e. the Dirichlet process when $\alpha=\gamma=0$. Between the two regimes, with $\mathbb{P}(r=1)=0.2$ resulting in $\mathbb{P}(r=1|\mathbf{Y})=0.5474$, the Bayes factor $B_{10}=\displaystyle\frac{\mathbb{P}(r=1|\mathbf{Y})}{\mathbb{P}(r=0|\mathbf{Y})}\left/\frac{\mathbb{P}(r=1)}{\mathbb{P}(r=0)}\right.=4.8368$, suggesting a slight preference for the finite regime, for this citation network.
Figure 4: Posterior histogram of $K_{n}$ and $k$, for the finite (solid) and infinite (stripe) regimes of the DAG-SBM, and model selection (crosshatch), for the SNA citation network. The right panel is for finite regime and model selection only as $k$ does not exist in the infinite regime. For readability, values of $k$ above $12$ are not shown.
@@
Figure 7: The posterior density of the positions of the vertices in $\boldsymbol{\sigma}$ for the infinite (left) and finite (right) regime of the DAG-SBM, for the SNA citation network. Each row represents the posterior density of a component of $\boldsymbol{\phi}$. The coloured dots are the mean positions.
@@
Next, we look at the posterior of $\boldsymbol{\sigma}$, or equivalently the positions of the vertices in the topological ordering, $\boldsymbol{\phi}$. The mixing of the MCMC is good, and the trace plots for some components of $\boldsymbol{\phi}$ are shown in the online Appendix D.2. The posterior density of each component of $\boldsymbol{\phi}$ is plotted as a row in Figure 7, with the rows themselves in an arbitrary topological ordering. As the citations, i.e. the edges in the DAG in general go from more recent works to older ones, the top (bottom) rows, which correspond to topologically earlier (later) vertices, are generally the more recent (older) articles in the network. The coloured part of each row represents the support of the posterior of the corresponding article, and in general is narrower towards the bottom. An interpretation is that generally chronologically older works are likely to cite between themselves and get cited by later ones, thus limiting their positions in $\boldsymbol{\sigma}$. Lastly, using Equation 8 and the SALSO algorithm (Dahl et al., 2022) to minimise the loss function, the point estimate $\hat{\mathbf{Z}}_{n}$ is obtained, with the co-clustering and adjacency matrices plotted in Figure 8.
@@ 3.3 Statistics citation network
We apply the model to a second citation network. The original data analysed by Ji and Jin (2016) contains 5722 citations between 3248 articles in the top statistics journals, from 2003 to the first half of 2012. We only consider the largest connected component, and remove one of the two edges in every pair of cyclic citations (there were only 9 pairs of these). Upon the data cleaning, we arrive at a citation network that is a DAG and contains 5563 citations between 2248 articles. We refer to this data as the statistics citation network hereafter.
The MCMC sampler was applied with $1.5\times 10^{4}$ iterations obtained after a burn-in period of $5.5\times 10^{4}$ and thinning of 10, with the same priors for $a$, $b$, $\alpha$, $\theta$, $k$ and $\gamma$ as those in Section 3.2. For model selection, even with $\mathbb{P}(r=0)=0.99$, the whole chain stays in the finite regime, i.e. $\mathbb{P}(r=1|\mathbf{Y})=1$. Therefore, we shall report the results under the finite regime only, as it is heavily preferred to the infinite counterpart for this network. The trace plots and posterior densities are provided in the online Appendix D.3. The posterior of $K_{n}$ ranges from 92 to 124, which seems reasonable with around 30 articles per cluster.
@@
Table 2: Titles of top articles according to topological ordering for selected groups.
@@
We report several groups of top articles according to median topological ordering for some large groups in Table 2. Specifically, for a selection of blocks, we report the top 5 articles in topological ordering within the block. From the article titles, groups seem to roughly correspond to various topics in statistics. For example, group 18 seems to cover false discovery rates, group 45 design and randomisation, group 2 wavelets, group 71 missing/perturbed data, group 25 functional linear regression, and group 65 spatial processes. However, note that the reported group memberships are just point estimates obtained from the MCMC draws, and therefore there might be some possibly large posterior uncertainty in the topological ordering for some of these articles.
Lastly, we report in Table 3 some top articles according to the median topological ordering, this time without the group memberships. In general, these articles have high number of citations. However, we also notice that some of these articles might not have a very high number of citations. This is due to the fact that these articles are cited by some other influential articles towards the end of the topological ordering.
Table 3: Titles of top 20 articles according to topological ordering.
@@ 4 Discussion
In this article, we proposed a Bayesian nonparametric SBM for DAGs. By conditioning on a latent topological ordering, the likelihood of the data (which is composed of directed edges) becomes that of an upper diagonal adjacency matrix. The topological ordering is treated as an unknown parameter, endowed with a prior and inferred a posteriori within the MCMC sampler, using a modified leap-and-shift proposal, previously used for the ranking in the Mallows model by Vitelli et al. (2018). The use of the PY process prior for the allocation vector $\mathbf{Z}_{n}$ allows the model to infer the number of groups $K_{n}$ from the data. Moreover, a model selection step for the two regimes of the PY process (one in which the total number of blocks is finite and unknown, and another one in which it increases with the sample size) is incorporated in the MCMC sampler.
The model can be generalised in different ways. For example, it can be extended by introducing covariate information, such as the keywords of each document or its publication year. This could be achieved by modelling the degree correction factors with a covariate dependent distribution. Also, in terms of parametrisation, the two regimes of the PY process could be unified so that $\gamma$ and $\alpha$ become one parameter that can take a value between -$\infty$ and $1$, with its posterior density directly implying which regime is preferred. The main obstacle to overcome here would be the sampling from the non-standard joint parameter space of $\theta$ and $\alpha$ across the two regimes. Another issue to be resolved is the inference of $k$, which is naturally highly correlated with $K_{n}$, under the finite regime. This is apparent in the parameter trace plots (in online Appendix D.3) for the statistics citation network, while for the SNA citation network the model selection improves the mixing of $k$ in the MCMC. Ideally $k$ is integrated out, but the computations required mean that this is feasible only under certain special cases. Such issue with $k$ remains to be resolved.
@@
Funding Acknowledgements
This research received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
Disclosure Statement
The authors report there are no competing interests to declare.
@@
Background materials of SBM and PY process, likelihood derivations, MCMC algorithms, and additional plots (appendices.pdf, pdf file).
@@
R-package “dagsbm” containing code to run the MCMC sampler for the DAG-SBM, and to simulate graphs from the model. The package also contains the dataset of SNA citation network presented in the article. (dagsbm_0.0.8.tar.gz, GNU zipped tar file)
@@
Script for running the MCMC sampler using the R package. (sim_dc.R, R script)
@@ Results:
Pre-generated results from running sim_dc.R, for the README and reproducibility checks. (results/*sim_dc_alpha=0.00_theta=1.00_n=250*.rds, Rds file)
@@
Instructions for reproducing the results in the paper using the R package and code. (README.Rmd, R Markdown script)